Morrow Company uses the high-low method to estimate the cost function. The information for 2020 is provided below:

                                                                               Machine-hours        Labor Costs

        Highest observation of cost driver                     900                       $36,000

        Lowest observation of cost driver                      100                       $14,000

What is the estimated total cost when 450 machine-hours are used?

A. 23,625

B.32,000

C. 18,000

D. 63,000
Answer #1

Variable cost per unit = (Highest activity cost - Lowest activity cost)/(Highest activity - Lowest activity)

= (36,000-14,000)/(900-100)

= 22,000/800

= $27.5 per machine hours

Fixed cost = Highest activity cost - Highest activity x Variable cost per hour

= 36,000 - 900 x 27.5

= 36,000-24,750

= $11,250

the estimated total cost when 450 machine-hours are used = Fixed cost + Variable cost per hour x Activity hour

= 11,250 + 27.5 x 450

= 11,250+12,375

= $23,625

Correct option is A.

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks